OpenStudy (mantar0078):

Write the augmented matrix, and use row reduction to solve. {2y=x+1 {3x-2=y Possible answers: A. (1, 1) [-1 2 : 1] [3 -1 : 2] B. (1, 2) [-1 2 : 1] [3 -1 : 2] C. (2, -1) [-1 1 : 2] [3 2 : -1] D. (-1, 1) [-1 1 : 2] [3 2 : -1

OpenStudy (unklerhaukus):

ax+by = e cx+dy = f [ a b ] [ x ] = [ e ] [ c d ] [ y ] = [ f ] [ a b : e ] [ c d : f ]

OpenStudy (mantar0078):

Is that how I will solve all equations like this one?

OpenStudy (unklerhaukus):

Yeah, first get them into the form ax + by = e cx + dy = f

OpenStudy (mantar0078):

So would I subtract x from the first one to get -x+2y=1 and how would I switch around the second one?

OpenStudy (unklerhaukus):

{2y = x+1 {3x-2 = y that's right for the first equation. 2y = x+1 (taking away x form both sides) -x+2y=1 now for the second 3x-2 = y (take away y from both sides) (and add 2 to both sides)

OpenStudy (mantar0078):

So I would get 3x-y=2?

OpenStudy (unklerhaukus):

Good! you now have {-x +2y = 1 {3x -y = 2

OpenStudy (unklerhaukus):

Which is of the form ax+by = e cx+dy = f We can express this system of equation as a matrix equation [ a b ] [ x ] = [ e ] [ c d ] [ y ] = [ f ]

OpenStudy (mantar0078):

So my answer would be [-1 2 : 1] [3 -1 : 2]

OpenStudy (unklerhaukus):

Yeah, that is the right augmented system

OpenStudy (unklerhaukus):

but you still have to solve it for x and y

OpenStudy (mantar0078):

How do I do that?

OpenStudy (unklerhaukus):

you have to reduce the argument system so something like [1 0 : g] [0 1 : h] which means the same thing as [ 1 0 ] [ x ] = [ g ] [ 0 1 ] [ y ] = [ h ] ie x = g and y = h

OpenStudy (unklerhaukus):

to reduce the augmented system you can take steps of either * multiplying entire rows by a constant * adding rows together

OpenStudy (unklerhaukus):

[-1 2 : 1] [3 -1 : 2] first step might be to change the first row by multiplying by -1

OpenStudy (unklerhaukus):

like this: R1 --> -R1 [1 -2 : -1] [3 -1 : 2] (now the element on the top left is a 1)

OpenStudy (mantar0078):

So my answer for x and y would be (1, 1)?

OpenStudy (unklerhaukus):

next step might be to change the second row by taking away three times the first row R2 --> R2 - 3R1

OpenStudy (unklerhaukus):

What do you get after this operation?

OpenStudy (mantar0078):

[0 -7 : -1]?

OpenStudy (unklerhaukus):

not quite [ 1, -2 : -1] [3-3(1), -1-3(-2) : 2-3(-1)] [ 1, -2 : -1] [3-3, -1+6 : 2+3] [ 1, -2 : -1] [ 0, -1+6 : 2+3]

OpenStudy (mantar0078):

[0 5 : 5]?

OpenStudy (unklerhaukus):

yeah we have [ 1, -2 : -1 ] [ 0, 5 : 5 ] (the bottom left element is now 0) now we should dived the second row by 5 R2 --> R2/5

OpenStudy (unklerhaukus):

[ 1, -2 : -1 ] [ 0, 5/5 : 5/5 ]

OpenStudy (mantar0078):

[0 1 : 1]

OpenStudy (unklerhaukus):

yeah we have [ 1, -2 : -1 ] [ 0, 1 : 1 ] (only one element left to go) Now, add two times the second row , to the first row R1 --> R1 + 2R2

OpenStudy (mantar0078):

[1 0 : 1]?

OpenStudy (unklerhaukus):

Great! We have [ 1, 0 : 1 ] [ 0, 1 : 1 ] aka [ 1 0 ] [ x ] = [ 1 ] [ 0 1 ] [ y ] = [ 1 ] which means 1x+0y = 1 0x+1y = 1 ie x = 1 y = 1

OpenStudy (unklerhaukus):

You might like to check that this solution solve both of the original equations.

OpenStudy (mantar0078):

Thank you so much for your help!